OpenStudy (prince1342):

Solve the following system of equations and show all work. y = x2 + 3 y = x + 5

OpenStudy (mecharv):

Is it x^2?

OpenStudy (prince1342):

yep!

OpenStudy (mecharv):

Okay lets use substitution method here... Substitute the second equation in first, So, x^2+3=x+5 x^2-x-2=0 Can you solve for x?

OpenStudy (prince1342):

2,-1

OpenStudy (mecharv):

Perfect Now substitute the values for x and find the corresponding y values :)

OpenStudy (prince1342):

Not sure how to do that hahaha

OpenStudy (mecharv):

I will help you :) Take any one equation... Let us take the linear one cause you know... quadratic is tougher than linear... y=x+5 so when x=2... y=2+5=7 So the solution pair is (2,7) Similarly find for x=-1 and thats your answer set :)

OpenStudy (prince1342):

2^2-1-2?

OpenStudy (mecharv):

Why did you go for the quadratic?? Use the same linear function no?

OpenStudy (mecharv):

y=x+5 what is y when x=-1?

OpenStudy (prince1342):

2 I think

OpenStudy (mecharv):

How ? O_o y=-1+5

OpenStudy (prince1342):

oh hahaha no it's 4

OpenStudy (mecharv):

Yes, so the answer set is (-1,4) and (2,7). Cheers Mate :)